Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (brand at Tue Jun 9 08:29:02 METDST 1998) Success

Torun (K. Borkowski at Tue Jun 9 08:46:08 METDST 1998) Success
No known problems.

Onsala (A. Polatidis at Wed Jun 10 10:54:43 METDST 1998) Success

Noto (C.Trigilio at Wed Jun 10 11:21:13 METDST 1998) Minor failures
Hardware problem to the recorder at the begin. The observation starts at 03:40.

Westerbork (Tony Foley at Wed Jun 10 11:41:52 METDST 1998) Minor failures
Strong EMI about 03:15UT. No other known problems

Jodrell (Alastair Gunn at Wed Jun 10 12:26:44 METDST 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Thu Jun 11 12:41:47 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the mai98 directory under the name ef005c.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 15:04:47 METDST 1998) Success